OpenStudy (anonymous):

I am completely confused , I just became a sophomore today, and I got Geometry, I do not understand anything!:(  I dont want to be behind. Here is one Q: The Fibonacci sequence consists of the patterns 1,1,2,3,5,8,13,...21,34... List the first rights terms of the sequence formed by the finding of the differences of the successive terms in the Fibonacci sequence 

OpenStudy (anonymous):

* eight terms

OpenStudy (zzr0ck3r):

2 = 1 + 1 3 = 2+1 5 = 3+2 8 = 5+3 do you see it?

OpenStudy (anonymous):

The fibonacci sequence follows a rule like the problem states. 1 and 1 is given as the first two terms. 1 + 1 = 2 1 + 2 = 3 2 + 3 = 5 3 + 5 = 8 5 + 8 = 13 8 + 13 = 21 13 + 21 = 34 21 + 34 = 55 34 + 55 = 89 55 + 89 = 144 etc...

OpenStudy (campbell_st):

the differences will be 1 - 1 = 0 2 - 1 = 1 3 - 2 = 1 5 - 3 = 2 8 - 5 = 3 13 - 8 = 5 21 - 13 = 8 34 - 21 = 13 etc hope this helps

OpenStudy (anonymous):

What are the 8 forms though?

OpenStudy (alexwee123):

did you also know that the fibonacci sequence can be found in pascals triangle too? :0

OpenStudy (zzr0ck3r):

It can also be found in the tool song lateralus

OpenStudy (anonymous):

@Itsgabiez The first 8 terms are 1, 1, 2, 3, 5, 8, 13, 21 These are the first eight terms. Can you see how the terms are found?

OpenStudy (alexwee123):

well i just thinks thats cool :/

OpenStudy (zzr0ck3r):

yeah for sure:)

OpenStudy (alexwee123):

@zzr0ck3r what's lateralus?

OpenStudy (zzr0ck3r):

a tool song

OpenStudy (zzr0ck3r):

TOOL is a band

OpenStudy (anonymous):

@Calmathlete is it by adding th first two terms then doing the same cycle? Like 1+1=2+3=5?

OpenStudy (alexwee123):

sounds cool might go listen to it :)

OpenStudy (anonymous):

Yes. The first two terms 1 and 1 are a given. 1 + 1 = 2 1 + 2 = 3 2 + 3 = 5 Yes. You take the result and add it to the previous term.

OpenStudy (anonymous):

Okay, thank ! Also I have another question, but I can't explain it on here because their diagrams... Like the Q says : draw the next figure in each sequence, like right triangle for example, why do I do for that?

OpenStudy (anonymous):

You can use the draw function. Or if the file is on the computer, just take a screenshot.

OpenStudy (anonymous):

& Sevenpeople meet and shake hands with one bother, using inductive reasoning, write a formula for the number of handshakes if the number of people is N :o that is absolutely confusing

OpenStudy (anonymous):

Yeah. I'm a bit confused about that to. Why don't you close this question and post an open one so that other people can help you?

OpenStudy (anonymous):

I took a screenshot, ok here is the link www.pic.twitter.com/wWtdpGuZ

OpenStudy (anonymous):

I'm using a mobile device

OpenStudy (anonymous):

It's a dead link.

OpenStudy (anonymous):

For 13, just draw a hexagon since the pattern is that one more side is being added. For 14, just draw a circle with 16 slices since the number of slices is being multiplied by 2.